The following sales and stock plans are available for the months of August through January.

                                                                                                Month Sales    BOM Stock                EOM Stock

August                                                                                    $ 77,678          $ 144,987                    $ 344,567

September                                                                             123,455             344,567                       456,986

October                                                                               99,678             456,986                       509,325

November                                                                          145,543             509,325                       602,345

December                                                                             245,778             602,345                       357,654

January                                                                              66,213             357,654                       251,345

What is the stock to sales ration for August?
Answer #1

Stock to Sales Ratio = Value of Average Stock in hand of a period / Value of Sold stock in that period

For August month BOM stock is = $144,987 and EOM Stock is =$ 344,567

Average Stock= (144,987 + 344,567) /2 = 489,554/2 = 244,777

and Sales of the month is = $ 77,678

Stock to sales ratio = 244,777 / 77,678 = 3.151 or 3.151:1

Stock to sales ratio indicates how much stocking is done by the business in comparison to sales, It is continuosly reviewed by businesses to avoid overstocking.

(Instead of Average stock in cases this ratio may also be calculated by taking stock at the beginning/end but in this case as there is a good gap in BOM and EOM stock levels, so application of Average stock is more appropriate.)